OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of one half. y - 6 = one half(x + 3) y = one halfx - 6 y + 3 = one half(x - 6) x - 2y = 12

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of one half. y - 6 = one half(x + 3) y = one halfx - 6 y + 3 = one half(x - 6) x - 2y = 12

OpenStudy (turingtest):

a line with slope \(m\) that passes through point \(a,b\) has equation\[y-b=m(x-a)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@TuringTest you still there

OpenStudy (anonymous):

i think the answer is either a or c

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

well given that I said that a line with slope \(m\) that passes through point \((a,b)\) has equation\[y-b=m(x-a)\]ask yourself, what is are \(a\) and \(b\) in your case?

OpenStudy (anonymous):

i have no clue to be honest

OpenStudy (turingtest):

what is the point the line passes through? it is given in the problem

OpenStudy (anonymous):

6,-3

OpenStudy (turingtest):

ok, so compare \((6,-3)\) and \((a,b)\) which number is \(a\) and which number is \(b\) ?

OpenStudy (anonymous):

6=a -3=b

OpenStudy (turingtest):

great! and remember that \(m\) is the slope what is the slope in your problem?

OpenStudy (anonymous):

ummm 1/2

OpenStudy (anonymous):

or one half

OpenStudy (turingtest):

correct, so \(m=\frac12,~a=6,~b=-3\) now just plug those numbers into\[y-b=m(x-a)\]

OpenStudy (anonymous):

y+3=1/2(x - 6)

OpenStudy (turingtest):

excellent! good job :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

wait i am done that is all have to do wow you are amazing do you mind helping me with some more

OpenStudy (anonymous):

i am going to close out of this though so i can give you another medal i will tag you

OpenStudy (anonymous):

@TuringTest

OpenStudy (turingtest):

sorry, I am multi-tasking you are welcome